OpenStudy (anonymous):

simplify 24x^3/ 8x^-2

OpenStudy (anonymous):

okay, \[\frac{24x^3}{8x^{-2}}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{x^{n}}{x^{m}}=x^{n-m}\]

OpenStudy (anonymous):

divide the coefficients

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{24}{8}=3\] \[\frac{x^{3}}{x^{-2}}=x^{3-(-2)}\]

OpenStudy (anonymous):

Not even going to say thanks?